### v2

# AlignSurvey: A Comprehensive Benchmark for Human Preferences Alignment in Social Surveys

Abstract
Understanding human attitudes, preferences, and behaviors through social surveys is essential for academic research and policymaking. Yet traditional surveys face persistent challenges, including fixed-question formats, high costs, limited adaptability, and difficulties ensuring cross-cultural equivalence. While recent studies explore large language models (LLMs) to simulate survey responses, most are limited to structured questions, overlook the entire survey process, and risks under-representing marginalized groups due to training data biases. We introduce AlignSurvey, the first benchmark that systematically replicates and evaluates the full social survey pipeline using LLMs. It defines four tasks aligned with key survey stages: social role modeling, semi-structured interview modeling, attitude stance modeling and survey response modeling. It also provides task-specific evaluation metrics to assess alignment fidelity, consistency, and fairness at both individual and group levels, with a focus on demographic diversity. To support AlignSurvey, we construct a multi-tiered dataset architecture: (i) the Social Foundation Corpus, a cross-national resource with 44K+ interview dialogues and 400K+ structured survey records; and (ii) a suite of Entire-Pipeline Survey Datasets, including the expert-annotated AlignSurvey-Expert (ASE) and two nationally representative surveys for cross-cultural evaluation. We release the SurveyLM family, obtained through two-stage fine-tuning of open-source LLMs, and offer reference models for evaluating domain-specific alignment. All datasets, models, and tools are available at github and huggingface to support transparent and socially responsible research.
Codehttps://github.com/PiLab-ZJU/AlignSurvey
Datasets & Modelshttps://huggingface.co/PiLabZJU
Extended Versionhttps://arxiv.org/abs/2511.07871
Introduction
Understanding human preferences, attitudes, and behaviors is central to both academic research and evidence-based policymaking. Social surveys, spanning qualitative interviews and quantitative questionnaires, have long served as a critical tool in this endeavor (wright2010survey; tourangeau2004survey), with an estimated \$35.1 billion spent annually on survey research worldwide (esomar2024global).
Despite this scale, traditional surveys face persistent challenges: fixed-question formats limit adaptability (heffetz2019difficulty); high costs often force sampling compromises, introducing biases (kalton2009methods); slow turnaround hinders responsiveness to emerging issues (moy2016problems; evans2018value; prosser2018twilight); and cross-cultural equivalence remains difficult despite translation efforts (tsai2025challenges).
Large language models (LLMs) offer a promising alternative. By learning from vast public corpora, LLMs can simulate human responses and reduce the burden of manual data collection (thapa2025large; mellon2024ais; zhang2025socioverse). However, their outputs often reflect the preferences of digitally active, well-educated users (giorgi2025human; abeliuk2025fairness), reinforcing representational bias (wang2025large; hu2025generative; hofmann2024ai) and marginalizing rural, low-income, or elderly populations. Group-level disparities are often underexplored, and many existing works lack systematic evaluation across demographic subgroups.
Figure 1: Overview of the AlignSurvey. AlignSurvey is a four-stage benchmark that mirrors the professional social survey process. The upper panel depicts a multi-tiered dataset: we pretrain on the Social Foundation Corpus for broad social science knowledge and fine-tune on the Entire-Pipeline Survey Dataset for the four survey stages: (1) Social Role Modeling; (2) Semi‑Structured Interview Modeling; (3) Attitude Stance Modeling; (4) Structured Response Modeling. AlignSurvey is the first to align LLMs across the entire social science survey, surpassing prior work limited to structured responses.
Existing benchmarks primarily focus on fixed-option quantitative tasks (ji2023beavertails; liu2025towards; zhou2025fair; lee2024can), overlooking the full pipelines of professional surveys, such as qualitative interviewing and context-aware reasoning. This critical gap calls for a comprehensive framework that aligns LLMs with human preferences across the entire survey process.
To address these gaps, we introduce AlignSurvey, the first benchmark designed to systematically replicate and evaluate the full pipeline of professional social surveys using LLMs. AlignSurvey mirrors four core stages of professional social surveys (ahmed2024mixed; fetters2013achieving) by mapping role exploration, qualitative survey, attitude mining and quantitative survey to corresponding modeling tasks: Social Role Modeling, Semi-structured Interview Modeling, Attitude Stance Modeling and Structured Response Modeling. We introduce task-specific evaluation metrics that enable alignment assessment across tasks at individual and group levels, with a particular focus on fairness and demographic diversity.
To support this pipeline, we construct a multi-tiered dataset architecture. The first component, the Social Foundation Corpus, is a cross-national resource comprising 44,000+ interview dialogues collected from publicly accessible video platforms and oral history books, and 400,000+ structured records from four authoritative surveys: ATP, ESS, CSS, and CGSS. These corpora provide foundational knowledge across diverse socio-cultural contexts, covering domains such as family dynamics, civic engagement, and inequality. We further introduce the Entire-Pipeline Survey Datasets for task-specific supervision. At the core is AlignSurvey-Expert (ASE), an expert-annotated dataset consisting of 161 semi-structured interviews and 1,679 questionnaires, comprising 2,500+ dialogues and 16,000+ responses, and enriched with detailed demographic metadata and annotated reasoning chains. To enable cross-national validation, we include two nationally representative datasets: the U.S.-based GSS and the China-based CHIP.
Benchmark Type Size Source Demographic Multi-Country Availability
Psychology-related
PhDGPT (de2024phdgpt) Quant 756K Synthetic Dataset
Psych-101 (binz2024centaur) Quant 10M Public Model + Dataset
Scenario (cui2025large) Quant // Public Dataset
Social Survey-related
OpinionQA (Santurkar2023WhoseOD) Quant 80K Public Dataset
Anthology (moon2024virtual) Quant 10K Public Dataset
SubPOP (suh2025language) Quant 73K Public Dataset
AlignSurvey (Ours) Qual + Quant 600K Expert + Public Model + Dataset
Table 1: Overview of alignment benchmark related to psychology and social surveys. “Quant” (quantitative) includes structured tasks with closed-form responses or ratings (e.g., survey choices or preference scores); “Qual” (qualitative) refers to expert-annotated open-ended responses, such as interviews and reasoning chains; and “Qual+Quant” covers both. “Source” indicates data origin: “Synthetic” is model-generated, “Public” is real-world open data, and “Expert” involves expert annotation.
Based on AlignSurvey, we fine-tune three representative open-source LLMs to develop the SurveyLM model family. The training follows a two-stage alignment strategy: we first adapt base model on the Social Foundation Corpus to equip it with general social knowledge and discourse patterns, then fine-tune them on each task using the Entire-Pipeline Survey Datasets. These models serve as strong reference baselines for evaluating alignment across survey tasks.
Experiments demonstrate that general-purpose LLMs fail to reliably reproduce survey outcomes, particularly for underrepresented groups. In contrast, SurveyLM models yield substantial improvements, 10–20% gains in demographic alignment and stance prediction, highlighting the need for domain-specific adaptation in social applications of LLMs.
Our contributions can be summarized as follows:
$\bullet$ We introduce AlignSurvey, the first benchmark that systematically replicates the full pipeline of professional social surveys using LLMs, including tasks for social role modeling, interview simulation, attitude modeling, and structured response prediction.
$\bullet$ We construct a multi-tier dataset architecture comprising (i) the Social Foundation Corpus (44K+ interviews, 400K+ survey records), and (ii) Entire-Pipeline Survey Datasets for full-pipeline alignment, enabling robust, cross-national evaluation and task-specific supervision.
$\bullet$ We release the SurveyLM model family, trained via a two-stage alignment strategy, as reference models for evaluating task-specific alignment across the survey pipeline.
$\bullet$ We contribute AlignSurvey-Expert (ASE), a multi-tiered dataset comprising expert-annotated interviews and thematic survey responses. It includes attitudinal stances, reasoning chains and demographic profiles, supporting supervision and evaluation across entire social survey process.
$\bullet$ All datasets, models, and evaluation code will be released at github and huggingface to support transparent, reproducible, and socially impactful research.
Related Work
Simulating Survey Respondents with LLMs.
To reduce the cost and rigidity of traditional survey methods (wright2010survey; heffetz2019difficulty; kalton2009methods), recent research explores using LLMs as virtual respondents (zhou2024exploring; zhu2025llm; wang2024human; li2024political; santurkar2023whose; abdurahman2024perils). Most approaches fall into prompt engineering, that simulates demographic variation via persona‐based instructions (santurkar2023whose; Hwang2023AligningLM; Simmons2022MoralML; Kim2024FewshotPO; Sun2024PersonaDBEL; moon2024virtual) and fine-tuning on real-world corpora to model individual or group preferences (Chu2023LanguageMT; He2024CommunityCrossInstructUI; Kwon2023EfficientMM; Zhao2023GroupPO; suh2025language). However, most of these efforts focus on isolated tasks and rarely address qualitative interviewing or full survey pipelines.
Benchmarks for Aligning LLMs with Social Survey Data.
While aligning LLMs with human preferences remains a core challenge (Kopf2023OpenAssistantC; Aroyo2023DICESDD; Lambert2024RewardBenchER; Ethayarajh2024KTOMA), benchmarks focused on social surveys are limited. Table 1 summarizes representative datasets across psychology and survey domains. Psychology-related resources have grown rapidly, often relying on synthetic scenarios to model cognition and responses (Binz2023TurningLL; binz2024centaur; Abdurahman2024PerilsAO; dominguez2024questioning). However, they mostly focus on individual cognition, lack real-world demographic variation, and do not model full survey workflows. Survey-focused benchmarks such as OpinionQA, Anthology, and SubPOP (Kirk2024ThePA; Santurkar2023WhoseOD; moon2024virtual; suh2025language) target structured, fixed-option questions, typically in U.S. contexts. They overlook qualitative interviewing, reasoning chains, and cross-cultural validity. None of these benchmarks replicates the entire process of professional social surveys or enables integrated evaluation across qualitative and quantitative components with expert annotations.
AlignSurvey
Design Principle
AlignSurvey is designed to systematically evaluate whether large language models can replicate the entire social survey process. As shown in Figure 1, the typical process of a professional social survey comprises four stages (ahmed2024mixed; fetters2013achieving): (1) Role Exploration: finding investigation targets, (2) Qualitative Survey: conducting semi-structured interviews, (3) Attitude Mining: extracting attitude stances and reasoning chains from dialogues, and (4) Quantitative Survey: collecting structured responses. AlignSurvey mirrors this process into four modeling and evaluation tasks: Social Role Modeling, Semi-structured Interview Modeling, Attitude Stance Modeling, and Structured Response Modeling. The Attitude and Response Modeling tasks include individual- and group-level subtasks to enable analysis of model alignment across multiple levels of social understanding. AlignSurvey provides stage-specific metrics to diagnose model strengths and limitations to support both rigorous benchmarking and responsible use in empirical social research.
Data Collection and Processing
AlignSurvey builds a multi-tiered dataset that combines large-scale public data and expert-curated resources to support comprehensive social contextual grounding and full pipeline alignment for LLM-based social surveys process.
Social Foundation Corpus.
This corpus provides foundational training to equip models with broad social knowledge and cultural patterns before task-specific alignment. It comprises two components:
Qualitative Corpus contains 44,021 structured interview dialogues from publicly accessible video platforms and books, spanning diverse national and cultural contexts. Dialogues are segmented into conversational turns, with a 5-turn sliding window used to predict the next response.
Quantitative Corpus comprises 411,174 records from four authoritative cross-national surveys: the American Trends Panel (ATP), European Social Survey (ESS), Chinese Social Survey (CSS), and Chinese General Social Survey (CGSS). These datasets cover topics such as public opinion, trust, inequality, and civic engagement. We selected five waves of ATP (40, 41, 54, 81, 103), Round 11 of ESS, and the latest releases of CSS and CGSS. To mitigate response biases (e.g., label-position bias (dominguez2024questioning)), option labels and choices were randomly shuffled.
Entire-Pipeline Survey Datasets.
As existing datasets seldom support full-pipeline alignment, we compile this suit of datasets, which comprises (1) AlignSurvey-Expert, an expert-annotated dataset, and (2) two national surveys from the U.S. and China.
AlignSurvey-Expert (ASE) is the core dataset for supervising and evaluating model performance across all four survey stages. It includes:
Qualitative Component includes 161 semi-structured interviews that comprise 2,500+ dialogues conducted by 15 social scientists, and each paired with rich demographic metadata (e.g., age, household size, occupation). The interviews are organized around eight core questions reflecting social perception. The first four gather general topic awareness, whereas the latter four target themes of service quality, social mobility, future expectations, and policy preferences. Each interview is divided into theme-specific dialogue segments, each comprising a sequence of utterances reflecting the respondent’s views. Each segment is annotated by six domain experts with an attitude label (positive, neutral, or negative) and a reasoning chain explaining the stance. See Appendix F for annotation details and reliability. These chains capture underlying logic by incorporating contextual factors such as life experience, group identity, and perceived fairness. Grounded in demographic profiles and dialogue content, they enable transparent, interpretable alignment evaluation in assessing subgroup-level fidelity, and serve as key supervision signals for training attitude models.
Quantitative Component includes 1,679 questionnaires that comprise 16,000+ responses designed around the same themes as the interviews. These were collected via an online platform, with rigorous screening applied to ensure data quality. Each question contains multiple labeled answer choices (e.g., A/B/C), which were randomly shuffled to mitigate position bias. The correct answer for each question is explicitly marked to support supervised learning and evaluation in structured response modeling.
Supplementary National Datasets. To support robustness checks and enhance cross-cultural generalization, we include two complementary national questionnaire datasets: the General Social Survey (GSS) from the U.S. and the China Household Income Project (CHIP). Both underwent the same preprocessing procedures as the questionnaire corpus of ASE, including label shuffling and standardized formatting, enabling consistent cross-dataset evaluation.
See Appendix A for detailed dataset information and Appendix G.1 for licensing details.
Task Definition
AlignSurvey defines four core tasks aligned with key stages of the social survey pipeline, each guided by tailored prompts and evaluation metrics to ensure faithful modeling and assessment.
ID Task Train Test
Task1 Social Role 8 712 2 880
Task2 Semi–structured Interview 1 904 632
Task3.1 Attitude Stance(Individual) 13 239 3 338
Task3.2 Attitude Stance(Group) 108 36
Task4.1 Structured Response(Individual) 54 180 12 047
Task4.2 Structured Response(Group) 46 815 8 674
Table 2: Details of tasks within AlignSurvey.
Task 1: Social Role Modeling.
This task evaluates whether an LLM can predict an interviewee’s demographic attribute (e.g., gender, education level) $c_{i}\in\boldsymbol{c}$ from a theme-specific dialogue segment $D^{(t)}$, where $t$ denotes the theme. The model prediction is: $\hat{c_{i}}=f_{LLM}(D^{(t)})$. Model performance is measured by accuracy: the proportion of predicted $\hat{c}_{i}$ matching the ground-truth $c_{i}$:
$Accuracy=\frac{1}{N}\sum_{i=1}^{N}\mathbf{I}[\hat{c}_{i}=c_{i}],$
where $\mathbf{I}[\cdot]$ is the indicator function.
Task 2: Semi-structured Interview Modeling.
This task evaluates whether the model can generate coherent interview responses conditioned on a full demographic profile $\boldsymbol{c}=(c_{1},c_{2},..,c_{i})$ and recent dialogue history $\mathcal{H}^{(t)}_{\tau}$, the last $\tau=5$ utterances of a dialogue $D^{(t)}$. The model prediction is: $\hat{d}^{(t)}_{\text{next}}=f_{LLM}(\boldsymbol{c},\mathcal{H}^{(t)}_{\tau})$. The generated response is evaluated along three dimensions: (i) Naturalness (fluency and topical coherence), (ii) Style Match (alignment with interview tone), and (iii) Consistency (logical and factual coherence).
Formally, let $\mathcal{J}$ be the set of evaluators  [^1] For Task 2 evaluation, to reduce bias and improve robustness, we use multiple strong LLMs (e.g., GPT-4o, DeepSeek-R1, Qwen-Max) as automated evaluators, with human validation confirming high agreement and consistency (see Appendix E.3)., and $s^{(d)}_{j,k}$ be the score assigned by evaluator $j\in\mathcal{J}$ on dimension $k$ for dialogue $d$. The final score is computed as:
$S^{(d)}_{k}=\frac{1}{|\mathcal{J}|}\sum_{j\in\mathcal{J}}s^{(d)}_{j,k}$.
Model Task 1 Task 2 Task 3
[…]
GPT-4o (Zero) 44.08% 35.28% 3.85% 3.45 2.85 2.89 38.12% 0.099 1.648
GPT-4o (Few) 40.59% 36.58% 17.43% 3.49 2.80 2.75 42.73% 0.103 1.180
Claude 3.7 Sonnet (Zero) 42.72% 34.17% 4.06% 3.41 2.81 2.85 28.12% 0.064 1.332
Claude 3.7 Sonnet (Few) 40.15% 33.18% 14.34% 3.48 2.83 2.90 31.48% 0.072 1.120
DeepSeek‑R1 (Zero) 38.52% 34.31% 6.56% 3.40 2.81 2.83 36.88% 0.075 1.356
DeepSeek‑R1 (Few) 36.36% 30.53% 18.27% 3.47 2.70 2.78 41.05% 0.081 1.100
R1‑Distill‑Qwen‑14B (Zero) 30.25% 27.08% 2.81% 2.78 2.20 2.36 38.75% 0.105 1.795
R1‑Distill‑Qwen‑14B (Few) 34.72% 31.18% 15.87% 2.85 2.25 2.30 33.81% 0.111 1.140
Qwen2.5‑72B (Zero) 37.83% 33.34% 3.12% 3.40 2.72 2.85 35.00% 0.102 1.653
Qwen2.5‑72B (Few) 38.65% 32.75% 16.90% 3.46 2.65 2.70 42.16% 0.113 1.050
Mistral‑7B‑v0.3 (Zero) 32.58% 35.14% 5.10% 3.21 2.57 2.44 38.44% 0.106 1.534
Mistral‑7B‑v0.3 (Few) 39.31% 32.73% 16.53% 3.28 2.50 2.40 39.77% 0.114 0.302
Meta‑Llama‑3.1‑8B (Zero) 37.00% 33.89% 8.33% 3.30 2.69 2.79 39.38% 0.107 1.521
Meta‑Llama‑3.1‑8B (Few) 35.42% 32.52% 16.02% 3.36 2.60 2.68 39.96% 0.119 0.447
Qwen2.5‑7B (Zero) 36.08% 33.89% 2.71% 3.39 2.67 2.79 36.88% 0.104 1.700
Qwen2.5‑7B (Few) 27.92% 37.83% 16.76% 3.44 2.60 2.85 44.69% 0.118 0.372
$SurveyLM_{\text{Mistral‑7B‑v0.3}}$ 48.24%∗∗ 52.23%∗∗ 47.08%∗∗ 2.61 2.90 2.90 57.13%∗∗ 0.131∗∗ 0.297∗∗
$SurveyLM_{\text{Meta‑Llama‑3.1‑8B}}$ 50.55%∗∗ 58.32%∗∗ 44.73%∗∗ 3.77 3.00 2.94 55.09%∗∗ 0.134∗∗ 0.458∗∗
$SurveyLM_{\text{Qwen2.5‑7B}}$ 54.65%∗∗ 57.23%∗∗ 48.71%∗∗ 3.98 2.96 3.01 56.83%∗∗ 0.128∗∗ 0.385∗∗
Table 3: Multi‑Task Evaluation Results. Even strong baselines such as GPT-4o still struggle, yielding low performance on fine-grained demographic prediction, interview style recognition, and stance detection. SurveyLM raises accuracy by about 10 to 15 percentage points and roughly halves Wasserstein distance. We conducted t-tests between each SurveyLM and its zero-shot model; $^{*}p<0.05$ and $^{**}p<0.01$. $\uparrow$ indicate that higher values are better, while $\downarrow$ indicate that lower values are better.
Task 3: Attitude Stance Modeling.
This task involves predicting both individual- and group-level attitude stances $\mathcal{A}$ and reasoning chains $\mathcal{R}$.
Individual-level. Given individual demographic profile $\boldsymbol{c}$ and theme $t$, the model predicts the attitude stance and reasoning chain: $(\hat{a}^{(t)},\hat{r}^{(t)})=f_{LLM}(\boldsymbol{c},t).$ Predictions are evaluated using macro-averaged accuracy, precision, recall, and F1 (for attitude stances), and ROUGE, Jaccard, and cosine similarity (for reasoning chains).
Group-level. For each question, the model produces individual stance predictions ${\hat{a}i^{(t)}}{i=1}^{N}$ for respondents belonging to a demographic group $t$. We convert these predictions into an empirical stance distribution by counting label frequencies: $\hat{\mathbf{p}}=\text{Aggregate}(\{\hat{a}_{i}^{(t)}\}),\quad\hat{\mathbf{p}}\in\Delta^{|\mathcal{A}|-1}.$ The predicted distribution is compared with the reference distribution $\mathbf{p}$ using the Wasserstein distance:
$W_{1}(\hat{\mathbf{p}},\mathbf{p})=\min_{\Gamma\geq 0,\Gamma\mathbf{1}=\hat{\mathbf{p}},\Gamma^{\top}\mathbf{1}=\mathbf{p}}\sum_{u,v\in\mathcal{A}}\Gamma_{uv}C_{uv},$
where $C_{uv}$ denotes the ground distance between stance labels $u,v\in\{\text{positive},\text{neutral},\text{negative}\}$.
To mitigate bias, we ground predictions in expert-annotated reasoning chains and training with demographic fine-grained profiles. This enables interpretable, group-aware alignment without relying on stereotypes.
Task 4: Survey Response Modeling.
This task predicts individual and group-level responses to questionnaire items.
Individual-level. Given a demographic profile $\boldsymbol{c}$ and question $Q_{i}$ with labeled options $(l_{i},o_{i})$, the model predicts the respondent’s selected answer $(\hat{l}_{i},\hat{o}_{i})=f_{LLM}(\boldsymbol{c},Q_{i})$, where $l_{i}$ denotes the label (e.g., A/B…) and $o_{i}$ denotes the content. Predictions are evaluated using accuracy, macro-averaged precision, recall, and F1, based on the ground-truth answers.
Group-level. Aggregating individual responses, the model generates a distribution over possible answers: $\hat{\mathbf{p}}_{q}=\text{Aggregate}(\{\hat{o}_{i}\}),\quad\hat{\mathbf{p}}_{q}\in\Delta^{|L|-1}.$ Alignment is measured via the Wasserstein distance (Similar to Task 3).
Table 2 summarizes the task definitions and dataset statistics. Further task-specific details and examples are provided in the Appendix D and E.
Constructing the SurveyLM Family.
We adopt a two-stage alignment strategy to build the SurveyLM model family. First, we fine-tune three open-source LLMs (Mistral 7B, LLaMA 3.1 8B, Qwen 2.5 7B) on the Social Foundation Corpus to equip them with general social concepts and culturally diverse discourse patterns. Then, we fine-tune the adapted models on the four AlignSurvey tasks using the Entire-Pipeline Survey Datasets, enabling task-specific alignment with both qualitative and quantitative objectives. The resulting models serve as strong reference baselines for domain-aligned social survey modeling and will be publicly released to support reproducible research.
Model ASE CHIP GSS
[…]
GPT-4o (Zero) 46.56% 22.05% 2.1336 53.43% 26.24% 1.9325 32.79% 24.41% 1.1130
GPT-4o (Few) 47.91% 23.10% 1.7895 54.32% 27.30% 1.6992 33.85% 25.10% 1.0284
Claude 3.7 Sonnet (Zero) 42.11% 18.44% 1.7789 47.21% 9.50% 1.8872 33.94% 25.93% 1.2248
Claude 3.7 Sonnet (Few) 42.95% 19.50% 1.6542 47.98% 9.95% 1.7318 34.55% 26.60% 1.1095
DeepSeek‑R1 (Zero) 47.21% 7.11% 1.9987 44.65% 4.52% 1.7789 34.54% 23.15% 1.3854
DeepSeek‑R1 (Few) 47.88% 7.65% 1.8527 45.27% 4.90% 1.6523 35.20% 23.70% 1.2432
R1‑Distill‑Qwen‑14B (Zero) 49.54% 25.90% 1.9295 50.62% 29.36% 1.9448 29.39% 11.57% 0.7920
R1‑Distill‑Qwen‑14B (Few) 50.29% 26.80% 1.7164 51.37% 30.67% 1.7104 30.20% 12.25% 0.7548
Qwen2.5‑72B (Zero) 42.20% 3.80% 1.8654 47.98% 13.29% 1.9918 34.50% 21.14% 1.2180
Qwen2.5‑72B (Few) 43.06% 4.05% 1.7348 48.66% 14.02% 1.7989 35.60% 22.05% 1.1112
Mistral‑7B‑v0.3 (Zero) 28.04% 1.10% 2.2118 23.60% 3.46% 1.9941 3.19% 0.20% 1.1479
Mistral‑7B‑v0.3 (Few) 30.47% 1.25% 2.0823 26.89% 3.70% 1.8876 10.30% 0.24% 1.0913
Meta‑Llama‑3.1‑8B (Zero) 14.84% 0.17% 2.1154 20.84% 2.40% 1.7748 1.61% 0.08% 1.5004
Meta‑Llama‑3.1‑8B (Few) 17.26% 0.19% 1.9451 22.46% 2.60% 1.6435 6.90% 0.12% 1.3817
Qwen2.5‑7B (Zero) 16.39% 0.12% 1.8479 51.95% 12.51% 1.9099 25.33% 4.30% 1.2411
Qwen2.5‑7B (Few) 18.46% 0.15% 1.7059 53.14% 13.15% 1.7441 26.85% 4.55% 1.1371
$SurveyLM_{\text{Mistral‑7B‑v0.3}}$ 54.59%∗∗ 41.34%∗∗ 0.0671∗∗ 65.56%∗∗ 35.35%∗∗ 0.0857∗∗ 51.06%∗∗ 35.73%∗∗ 0.2097∗∗
$SurveyLM_{\text{Meta‑Llama‑3.1‑8B}}$ 56.11%∗∗ 41.47%∗∗ 0.0675∗∗ 64.89%∗∗ 32.99%∗∗ 0.0845∗∗ 51.44%∗∗ 33.84%∗∗ 0.2137∗∗
$SurveyLM_{\text{Qwen2.5‑7B}}$ 55.47%∗∗ 41.79%∗∗ 0.0541∗∗ 67.17%∗∗ 34.94%∗∗ 0.0686∗∗ 51.52%∗∗ 36.22%∗∗ 0.1849∗∗
Table 4: Survey Response Modeling (Task4) Results. In zero- and few-shot settings, general-purpose LLMs give low-fidelity answers. On the ASE, CHIP, and GSS datasets, SurveyLM achieves high accuracy and F1 scores while yielding substantially smaller Wasserstein distances. The large F1 jump shows SurveyLM no longer defaults to midpoint guesses but distinguishes the full spectrum, reducing central-tendency bias. Two-sided t-tests vs. zero-shot baselines: $^{*}p<0.05$, $^{**}p<0.01$. $\uparrow$ indicate that higher values are better, while $\downarrow$ indicate that lower values are better.
Experiments & Discussion
Setup and Implementation Details
We evaluate models under three settings: zero-shot, few-shot (with three in-context examples), and supervised fine-tuning (SFT). SFT corresponds to the training procedure used to construct the SurveyLM family, which adapts LLMs through corpus-level and task-specific supervision. SFT first trains on the Social Foundation Corpus (1 epoch), followed by each task dataset (3 epochs). Prompts exceeding the model’s context window are truncated by removing the middle segment, retaining the task instruction and query. Prompt templates are provided in Appendix D and E.
[…]
Evaluated Models
[…]
Experimental Results & Analysis
Figure 2: A radar chart showing the overall performance on the tasks. Each axis represents the task score, with larger values indicating better performance. We used the inverse of the WD. Each broken line corresponds to a model, with larger enclosed areas indicating better overall performance. Overall, SurveyLM series performed exceptionally well.
Table 3 and Table 4 report aggregated results for Tasks 1–3 and Task 4. For Task 1, we divided accuracy into three categories which specific classification can be found in Appendix B.1. Figure 2 complements these with a radar chart overview of model performance across tasks. Detailed per-label scores and metrics are provided in Appendix B and C.
Task 1: Social Role Modeling.
SurveyLM models significantly outperform all baselines in predicting fine-grained demographic attributes. While top-tier models like GPT-4o perform decently on binary traits (e.g., gender), they struggle with complex categories such as family structure and social roles. Few-shot prompting yields limited gains. SurveyLM models improve accuracy by over 10 points across attributes, with the strongest gains in familial and social context fields. Notably, a fine-tuned 7B model (Qwen) surpasses its 72B counterpart, highlighting the impact of domain alignment over model scale.
Task 2: Semi-structure Interview Modeling.
General-purpose models (e.g., GPT-4o, DeepSeek-R1) reach moderate naturalness but lack stylistic consistency and fail to match the interview style. SurveyLM models, with two-stage alignment, yield more fluent, coherent, and contextually appropriate responses, demonstrating the importance of fine-tuning for qualitative generation.
Task3. Attitude Stance Modeling.
General-purpose models tend to default to majority stances with generic reasoning, showing 28–45% accuracy and limited few-shot improvement. They perform poorly on group-level metrics Wasserstein Distance. SurveyLM models exceed 55% accuracy and generate higher-quality explanations (Jaccard higher), confirming the value of alignment for accurate, interpretable attitude inference in public opinion modeling.
Task4. Survey Response Modeling.
In both zero- and few-shot settings, general-purpose models yield low-fidelity outputs (accuracy <<50%, low F1). SurveyLM achieves high accuracy and F1 scores while yielding substantially smaller Wasserstein distances on ASE, CHIP, and GSS. The pronounced jump in F1 shows that the model no longer “plays it safe” by predicting the midpoint response for everyone. It learns to recognize and output the full spectrum of options. This balanced improvement mitigates the central-tendency bias that questionnaires often suffer from. These results demonstrate the strength of supervised alignment for structured response prediction.
Ablation Analysis.
To assess the impact of each alignment stage, we conduct ablation tests on three backbones (Mistral-7B, LLaMA-3.1-8B, Qwen2.5-7B), removing either Stage I (foundation adaptation, w/o Foundation) or Stage II (task-specific tuning, w/o Task SFT), keeping other conditions fixed.
Model T1 T2 T3 T4(ASE)
Acc Avg Acc WD Acc WD
$SurveyLM_{\text{Mistral}}$ 0.47 2.80 0.57 0.297 0.55 0.067
   w/o Foundation 0.46 3.10 0.56 0.303 0.53 0.077
   w/o Task SFT 0.38 2.96 0.44 0.313 0.36 2.182
$SurveyLM_{\text{Llama}}$ 0.45 3.24 0.55 0.458 0.56 0.068
   w/o Foundation 0.44 3.07 0.54 0.463 0.55 0.085
   w/o Task SFT 0.39 3.01 0.36 0.484 0.34 2.098
$SurveyLM_{\text{Qwen}}$ 0.49 3.32 0.57 0.385 0.56 0.054
   w/o Foundation 0.47 3.12 0.56 0.389 0.55 0.889
   w/o Task SFT 0.35 2.79 0.47 0.397 0.36 2.002
Table 5: Ablation results on Tasks 1–4. T1–T4 correspond to Tasks 1–4, respectively. For Tasks 1 and 2 we report accuracy and the average score; for Task 4 we report results on the ASE split only. Detailed results appear in Appendix C.1.
As shown in Table 5, removing either stage degrades performance. Omitting foundation adaptation yields moderate drops, while skipping task-specific tuning causes severe degradation. Full SurveyLM models outperform ablated variants, confirming that foundation-level pretraining and task-specific supervision are complementary. This highlights the effectiveness of our two-stage alignment strategy in adapting general-purpose LLMs to social survey tasks.
Figure 3: Multi‑Demographic Accuracy Comparison. Circle size represents the accuracy gap between Qwen-7B and its SurveyLM. Gains are especially pronounced for under‑represented groups (rural residents, aged 76+, self‑employed, low‑ and middle‑ income groups).
Equity-Oriented Gains.
Figure 3 shows Task 3 accuracy across demographic attributes using the Qwen2.5-7B. SurveyLM consistently shifts scores toward higher accuracy, while baselines often remain near or below midpoint. Notably, SurveyLM significantly boosts accuracy for underrepresented groups, including rural, elderly (76+), self-employed, and low-/middle-income individuals, reducing disparities with advantaged counterparts (urban, college-educated, high-income) and enhancing demographic equity. Full comparisons across models are in the Appendix C.2.
Social Impact and Alignment Implications
Our experiments show that socially grounded alignment enhances both performance and fairness. AlignSurvey and SurveyLM help recover signals from marginalized groups, reducing reliance on dominant narratives and supporting more inclusive policy diagnostics, welfare targeting, and equitable decision-making. More broadly, this work demonstrates how targeted alignment can bridge LLMs with real-world societal needs,especially in governance, auditing, and digital public services.
Conclusion & Future Work
This paper presents AlignSurvey, the first benchmark to systematically replicate the full pipeline of social surveys using LLMs. By integrating Social Foundation Corpus and Entire-Pipeline Survey Datasets, it enables comprehensive evaluation across demographic modeling, qualitative interaction, attitude inference, and structured response prediction. Experiments demonstrate alignment on AlignSurvey can recover signals from underrepresented groups, reducing bias and supporting more inclusive, policy-relevant modeling.
Future directions include iterative improvement through human-in-the-loop feedback and expanded coverage across diverse cultural settings for broader applicability.
Ethical Statement
We take ethics seriously. Our dataset is sourced from publicly accessible content, including video-platform APIs, research-cleared books, and licensed surveys. Personally identifiable information is removed or anonymized before training, and only privacy-preserving, non-harmful content is released. Raw audio, video, or verbatim transcripts are never shared.
We implement multiple safeguards to prevent misuse, including controlled data access, content filtering, and usage guidance tailored to responsible LLM-based survey applications. Data is used solely for non-commercial academic research, and compliance with privacy and copyright regulations is continuously monitored. See Appendix G for further details.
Acknowledgments
This research was supported by the National Natural Science Foundation of China (Grant Nos. 72574198, 72434004), the Key Project of Humanities and Social Sciences of the Ministry of Education of China (Grant No. 2023JZDZ038), and the National Social Science Fund of China (Grant No. 23BZZ088). We thank all collaborators contributing to data collection, annotations and valuable feedback.
A. Dataset Documentation
A.1 Social Foundation Corpus
The Social Foundation Corpus serves as the broad‐coverage pre-training layer for AlignSurvey. It consists of two coordinated components:
  1. 1.
    Qualitative component: 44,021 fully anonymised interview dialogues curated from public video platforms and published academic sources related to oral history. Dialogues longer than 2,000 tokens were centre–truncated and a five-turn sliding context window was generated for next-utterance prediction.
  2. 2.
    Quantitative component: 411,174 structured survey records drawn from four flagship probability surveys (ATP, ESS, CSS, CGSS), covering trust in institutions, inequality, religion, media use, family dynamics and civic engagement. Response options were randomly re-labelled to minimise order bias.
Details of each component are provided below.
Data Item Size
American Trends Panel(ATP) Quan. 106,028
European Social Survey(ESS) Quan. 98,303
Chinese Social Survey(CSS) Quan. 104,048
Chinese General Social Survey(CGSS) Quan. 102,795
Public Accessible Video Platforms Qual. 40,077
Textual Excerpts from Academic Sources Qual. 3,944
Table 6: Origin, data modality (Quan. = quantitative survey items; Qual. = qualitative materials), and item count of the foundational datasets used in this study.
American Trends Panel (ATP):
Pew Research Center’s ATP is the only fully probability-based online panel that tracks the same U.S. adults over time. Households without internet receive a tablet and mobile data plan, ensuring true national coverage. We selected five waves with strong economic-attitudes overlap with our benchmark tasks. These waves were chosen based on their relevance to key social survey themes used in dataset construction. The selected waves are listed in Table 7.
Wave Wave Top Population
40 Trust, facts, and democracy 10,618
41 Views of America in 2050 2,526
54 Economic inequality 6,878
81 Economics, pandemic financial outlook 10,334
103 Economic well-being 9,388
Table 7: ATP Survey Waves. Wave number, thematic focus, and sample size of the five American Trends Panel waves included in the quantitative tier of the Social Foundational Corpus.
European Social Survey (ESS):
Since 2002 the ESS has fielded biennial, face-to-face probability surveys that apply identical sampling rules, translated instruments and centralised data processing across participating nations. Rigorous design—random dwelling selection, strict interviewer monitoring and double back-translation—makes the ESS a gold standard for cross-national comparisons of attitudes, beliefs and behavioural patterns. We selected Round 11 of the ESS for dataset construction. It is the most recent publicly available wave, covering 28 countries, and includes a rotating module on Justice & Fairness in addition to the core questionnaire battery.
Chinese Social Survey (CSS):
Launched by the Institute of Sociology, Chinese Academy of Social Sciences in 2005, the CSS conducts biennial multi-stage, stratified probability interviews covering all 31 provincial-level units. By retaining a core module and rotating topical modules, CSS yields both time-series indicators of social change and flexibility to explore emerging issues in a rapidly transforming society. We selected the latest data from 2021 to build the dataset. Wave 2021 emphasised social mobility, government service quality and subjective class identity.
Chinese General Social Survey (CGSS):
Jointly organized by Renmin University of China and other institutions since 2003, the Chinese General Social Survey (CGSS) is the longest-running, nationally representative repeated cross-sectional survey of Chinese adults aged 18 and above. Its questionnaire combines fixed trend items with rotating modules covering topics such as social mobility, neighborhood life, political culture, and health. We selected the most recent wave from 2023 for dataset construction.
Publicly Accessible Video Platforms:
To simulate the conversational structure and thematic richness of in-person qualitative fieldwork, we collected 388 long-form interview videos on social issues from publicly accessible video platforms. Each recording was segmented into speaker turns (diarized), transcribed, and filtered for offensive language. All personal identifiers were anonymized prior to analysis. A full description of these preprocessing steps is provided in Appendix A.3.
Textual Excerpts from Academic Sources:
We included textual materials from published works on qualitative social surveys and oral history, extracting interviewer–respondent exchanges from the text. These materials were processed using the same pipeline as the video interviews, including speaker segmentation, transcription, and anonymization to ensure appropriateness for academic use.
Data Item
AlignSurvey-Expert(ASE) Qual.+ Quan.
General Social Survey(GSS) Quan.
China Household Income Project(CHIP) Quan.
Table 8: Entire-Pipeline Survey Datasets: overview of the three corpora that supervise the full modelling pipeline, with their respective data modalities (Qual. = qualitative; Quan. = quantitative).
A.2 Entire-Pipeline Survey Datasets
The Entire-Pipeline Survey Datasets supply explicit supervision for each of the four modelling stages—Social-Role Modelling, Semi-Structured-Interview Modelling, Attitude–Stance Modelling, and Survey-Response Modelling. An overview of all component corpora is given in Table 8, while corpus-level statistics for the Attitude, Stance, and Engagement (ASE) corpus appear in Table 3 of the main paper. The demographic profile collected for every respondent is summarised in Table 11.
Qualitative component.
The qualitative part of ASE comprises semi-structured interviews that cover eight broad themes (Table 10).
Gender Count Percent
Female 3239  53.6%
Male 2804  46.4%
Age
18–25 1945  32.2%
26–35 2151  35.6%
36–45 1103  18.3%
46–55 619  10.2%
56–65 192  3.2%
66–75 21  0.3%
$\geq 76$ 12  0.2%
Region
Urban 3300  54.6%
Rural 2743  45.4%
Occupation
Sales 1724  28.5%
Professional 979  16.2%
Other 968  16.0%
Student 871  14.4%
Government Employee 686  11.4%
Manual Labor 482  8.0%
Retired 160  2.6%
Business Owner 151  2.5%
Self Employed 10  0.2%
Company Employee 8  0.1%
Unemployed 4  0.1%
Education
Bachelor’s Degree 3365  55.7%
Associate Degree 1177  19.4%
Graduate Degree 800  13.3%
High School 556  9.2%
Junior High School 145  2.4%
Home Ownership
Yes 3112  51.5%
No 2931  48.5%
Car Ownership
Yes 3180  52.6%
No 2863  47.4%
Household Size
0 1  0.0%
1 62  1.0%
2 669  11.1%
3 2219  36.7%
4 1453  24.0%
5 959  15.9%
6 427  7.1%
7 126  2.1%
$\geq 8$ 127  2.1%
Income
110k–200k 1985  32.8%
210k–500k 1748  28.9%
60k–100k 1179  19.5%
$\leq 50k$ 647  10.7%
[…]
510k–600k 221  3.7%
Stress Level
Moderate Stress 2695  44.6%
Mild Stress 1898  31.4%
High Stress 911  15.1%
No Stress 351  5.8%
Very High Stress 188  3.1%
Table 9: Task4 Demographic Distribution with Visual Percent Bars
Quantitative component.
The quantitative part of ASE is a county-level development-and-governance questionnaire. Its hierarchical theme structure—parent topics and corresponding survey questions—is shown in Table 12.
ID Theme
1 Policy understanding
2 Policy details known
3 Policy awareness
4 Income level awareness
5 Public service evaluation
6 Social class attitude
7 Upward mobility attitude
8 Future policy expectations
Table 10: Interview themes in the ASE qualitative dataset.
Attribute Option Value
Gender Male; Female
Age 18–25 years; 26–35 years; 36–45 years; 46–55 years; 56–65 years; 66–75 years; 76 years and above
Region Urban; Rural
Education Primary School; Junior High School; High School; Associate Degree; Bachelor’s Degree; Graduate Degree
Home Ownership Yes; No
Car Ownership Yes; No
Household Income Below ¥50,000; ¥50,000–200,000; ¥200,000–500,000; ¥500,000–1,000,000; Above ¥1,000,000
Stress Level No Stress; Mild Stress; Moderate Stress; High Stress; Very High Stress
Occupation Student; Retired; Company Employee; Government Employee; Private Business Owner; Self-employed Merchant; Other
Registered Residence Geographical location
Current Residence Geographical location
Household Size Integer, e.g., 4
Family Relations e.g., spouse, children
Main Monthly Expenses e.g., mortgage, education, healthcare
Table 11: Demographic attributes collected for participants in the ASE.
ID Survey Question
Economic Development
1.1 Perception that the county’s economy is developing rapidly
1.2 Perception that the county’s economic‐development model is sustainable
1.3 Improvement in local job opportunities and salary levels
1.4 Government effort to promote high-quality economic development
Social Gaps and Fairness
2.1 Evaluation of the income gap
2.2 Evaluation of the urban–rural gap
2.3 Evaluation of the regional gap
2.4 Perception that local society is fair
Public-Service Satisfaction
3.1 Satisfaction with public education services
3.2 Satisfaction with public medical services
3.3 Satisfaction with public elderly-care services
3.4 Satisfaction with public childcare services
3.5 Satisfaction with public employment services
3.6 Satisfaction with public housing services
3.7 Satisfaction with public assistance services
3.8 Satisfaction with public transportation services
Governance and Participation
4.1 Perception that the social environment is safe and stable
4.2 Perception that the county’s governance level has improved
4.3 Ability to participate in community or village affairs
4.4 Satisfaction with the current government’s development performance
4.5 Confidence in the government’s future development
Table 12: Hierarchical topic structure of the ASE quantitative questionnaire: parent themes and their child survey questions
A.3 Pre-processing and Splits
All interview recordings in the qualitative datasets are first transcribed with an automatic speech-recognition system; subsequently, human editors sample and review 2 % of the material to bring the transcripts to near-perfect orthographic accuracy, intervening in disfluencies only when these obstruct comprehension. After transcription, we run a conservative, rule-based anonymiser that leverages GPT-4o: every overt personal identifier—names, institutions, addresses, and similar noun phrases—is replaced with a unique placeholder (e.g., “I, <NAME_17C><NAME\_17C>, was admitted to Harvard University in 1998”). Each interview alternates between interviewer (A) and interviewee (B), though turn lengths vary. We tokenise sessions into speaker-aligned turns and apply a five-turn sliding window that always begins with interviewer A and advances by one turn at a time (e.g., for A B A B A B A, window 1 spans turns 1–5, window 2 spans turns 3–7, etc.). Windows are discarded if their concatenated text, after trimming whitespace, contains fewer than 20 alphabetic tokens or if anonymisation has redacted more than 30 % of the tokens.
In the quantitative datasets, every single-choice answer is shuffled at load time, as described in the main paper. To standardise rating scales, Likert inventories with an odd number of anchors are remapped onto a canonical three-class backbone: edge bins are collapsed in proportion to each item’s empirical distribution so that the resulting histogram is as balanced as possible. Scales with an even number of anchors, which lack a neutral midpoint by design, are instead projected onto a four-class backbone. To prevent the model from memorising positional biases, we also shuffle the letters associated with each option as well as the option texts themselves.
For the Social Foundation Corpus, all datasets are ingested to fine-tune the backbone model via supervised instruction (SFT), thereby imparting foundational knowledge before evaluation. Test splits of the Entire-Pipeline Survey Datasets are designed to probe generalisation across both demographic profiles and thematic domains. Each respondent receives a fingerprint composed of six categorical pillars—gender, age, region status, highest educational level, household-income bracket, and occupation category. Respondents whose fingerprint never appears in the provisional training pool are assigned to the hold-out group, and they constitute exactly half of the final test set; the remaining half consists of respondents who do occur in training but answer different questions or segments, enforcing cross-topic rather than cold-start extrapolation. Concretely, the ASE split is reported in Table 3 of the main paper; the GSS split contains 18 922 training and 4 730 test instances, while the CHIP split comprises 11 718 training and 3 129 test instances.
B. Comprehensive Experimental Results
B.1 Task1 Social Role Modeling
Table 13 presents detailed results for Task 1: Social Role Modeling. This task aims to evaluate the capability of various large language models (LLMs) to infer demographic and socioeconomic attributes based solely on provided dialogues.
[…]
Few-shot prompts improve the performance of general models, such as GPT-4o (from 0.00% to 23.75% in ”Relationship”). However, even with few-shot learning, these models remain significantly below SurveyLM performance. Models with smaller scales generally show weaker performance, particularly in zero-shot settings, underscoring the importance of specialized fine-tuning for accurate demographic inference.
Model Basic INFO Social INFO Family INFO
Gender Age Area Edu. Employ. Stress Home Own. Veh. Own. Income Rel. HH Size Expenses
GPT-4o (Zero-shot) 50.00% 37.08% 53.75% 34.17% 45.42% 9.58% 47.92% 48.33% 11.67% 0.00% 3.75% 0.00%
GPT-4o (Few-shot) 52.25% 40.25% 39.50% 34.75% 36.20% 11.80% 48.10% 49.85% 19.22% 23.75% 2.50% 24.25%
Claude (Zero-shot) 51.92% 35.00% 57.92% 23.75% 45.00% 12.08% 47.50% 42.92% 7.50% 0.00% 7.92% 0.83%
Claude (Few-shot) 51.10% 40.10% 37.05% 39.10% 33.40% 11.40% 44.35% 43.80% 11.13% 21.25% 0.40% 24.60%
DeepSeek−R1 (Zero-shot) 51.33% 34.58% 55.83% 12.50% 38.33% 1.67% 52.08% 49.17% 8.75% 0.00% 15.83% 1.67%
DeepSeek−R1 (Few-shot) 51.05% 40.00% 42.60% 31.20% 16.95% 2.45% 54.55% 34.60% 11.67% 20.60% 17.65% 23.15%
R1−Distill−Qwen−14B (Zero-shot) 38.75% 21.67% 55.00% 14.17% 21.67% 0.00% 50.42% 30.83% 0.00% 0.00% 11.25% 0.00%
R1−Distill−Qwen−14B (Few-shot) 40.60% 43.60% 42.25% 24.05% 23.10% 2.55% 52.45% 38.55% 8.11% 20.30% 12.15% 22.90%
Qwen2.5−72B (Zero-shot) 48.20% 43.95% 50.51% 17.27% 29.24% 1.23% 48.77% 50.00% 4.27% 0.00% 8.22% 0.00%
Qwen2.5−72B (Few-shot) 46.05% 43.30% 47.55% 23.25% 33.10% 1.50% 60.95% 35.80% 9.67% 20.35% 14.60% 23.00%
Meta−Llama−3_1−8B (Zero-shot) 47.50% 40.42% 58.33% 12.50% 26.25% 0.00% 50.83% 50.83% 6.67% 0.00% 26.67% 0.00%
Meta−Llama−3_1−8B (Few-shot) 49.05% 40.05% 41.55% 23.10% 23.35% 4.40% 47.10% 46.05% 9.11% 20.45% 11.55% 22.95%
Mistral−7B−v0.3 (Zero-shot) 39.17% 35.00% 48.75% 4.58% 35.42% 3.33% 51.67% 50.42% 19.17% 0.00% 0.83% 0.42%
Mistral−7B−v0.3 (Few-shot) 22.80% 43.10% 43.35% 55.05% 32.25% 6.35% 52.60% 39.25% 13.33% 20.40% 9.45% 22.95%
Qwen2.5−7B (Zero-shot) 46.67% 41.67% 49.58% 15.83% 26.67% 0.83% 50.83% 50.00% 4.17% 0.00% 6.67% 0.00%
Qwen2.5−7B (Few-shot) 48.55% 43.45% 45.10% 17.25% 35.25% 1.65% 60.10% 51.75% 9.83% 20.55% 13.70% 22.95%
$SurveyLM_{\text{Mistral-7B-v0.3}}$ 54.31% 45.87% 59.67% 37.90% 43.47% 35.44% 60.54% 60.71% 33.18% 85.88% 39.99% 29.28%
$SurveyLM_{\text{Meta-Llama-3\_1-8B}}$ 55.14% 47.03% 62.05% 41.02% 47.49% 44.02% 65.92% 65.01% 31.29% 80.06% 34.80% 32.78%
$SurveyLM_{\text{Qwen2.5-7B}}$ 67.81% 50.62% 60.58% 41.83% 52.43% 36.55% 68.25% 66.88% 40.42% 86.42% 34.04% 33.95%
Table 13: Task1 Social Role Modeling Comprehensive Results.
B.2 Task2 Semi-Structured Interview Modeling
Table 14 presents evaluation scores on naturalness, style matching, and consistency of generated interview responses.
[…]
Model Naturalness $\uparrow$ Style Match $\uparrow$ Consistency $\uparrow$
GPT-4o (Zero-shot) 3.45 2.85 2.89
GPT-4o (Few-shot) 3.49 2.80 2.75
Claude 3.7 Sunnet (Zero-shot) 3.41 2.81 2.85
Claude 3.7 Sunnet (Few-shot) 3.48 2.83 2.90
DeepSeek-R1 (Zero-shot) 3.40 2.81 2.83
DeepSeek-R1 (Few-shot) 3.47 2.70 2.78
R1-Distill-Qwen-14B (Zero-shot) 2.78 2.20 2.36
R1-Distill-Qwen-14B (Few-shot) 2.85 2.25 2.30
Qwen2.5-72B (Zero-shot) 3.40 2.72 2.85
Qwen2.5-72B (Few-shot) 3.46 2.65 2.70
Meta-Llama-3.1-8B (Zero-shot) 3.30 2.69 2.79
Meta-Llama-3.1-8B (Few-shot) 3.36 2.60 2.68
Mistral-7B-v0.3 (Zero-shot) 3.21 2.57 2.44
Mistral-7B-v0.3 (Few-shot) 3.28 2.50 2.40
Qwen2.5-7B (Zero-shot) 3.39 2.67 2.79
Qwen2.5-7B (Few-shot) 3.44 2.60 2.85
$SurveyLM_{\text{Mistral-7B-v0.3}}$ 2.61 2.90 2.90
$SurveyLM_{\text{Meta-Llama-3 1-8B}}$ 3.77 3.00 2.94
$SurveyLM_{\text{Qwen2.5-7B}}$ 3.98 2.96 3.01
Table 14: Task2 Semi-Structured Interview Modeling Result.
B.3 Task3 Attitude Stance Modeling
Table 15 outlines detailed metrics evaluating attitude stance prediction.
[…]
Model ACC $\uparrow$ PRE $\uparrow$ RECALL $\uparrow$ F1 $\uparrow$ ROUGE $\uparrow$ Jaccard $\uparrow$ Cosine $\uparrow$ WD $\downarrow$
GPT-4o (Zero-shot) 38.12% 32.07% 32.41% 29.35% 9.83 0.0987 0.0125 1.6479
GPT-4o (Few-shot) 42.73% 35.46% 34.04% 32.78% 9.10 0.1031 0.0137 1.1800
Claude 3.7 Sonnet (Zero-shot) 28.12% 20.72% 17.71% 17.35% 8.16 0.0637 0.0002 1.3319
Claude 3.7 Sonnet (Few-shot) 31.48% 23.63% 20.97% 19.89% 8.10 0.0715 0.0006 1.1200
DeepSeek-R1 (Zero-shot) 36.88% 35.05% 33.50% 32.06% 8.19 0.0745 0.0026 1.3561
DeepSeek-R1 (Few-shot) 41.05% 36.27% 34.66% 33.86% 8.12 0.0810 0.0034 1.1000
R1-Distill-Qwen-14B (Zero-shot) 38.75% 35.09% 34.93% 34.36% 8.44 0.1045 0.0058 1.7952
R1-Distill-Qwen-14B (Few-shot) 33.81% 29.11% 28.46% 28.02% 8.20 0.1111 0.0076 1.1400
Qwen2.5-72B (Zero-shot) 35.00% 32.51% 31.04% 31.05% 9.95 0.1015 0.0097 1.6533
Qwen2.5-72B (Few-shot) 42.16% 38.92% 37.42% 36.90% 8.35 0.1126 0.0141 1.0500
Mistral-7B-v0.3 (Zero-shot) 38.44% 35.11% 33.98% 33.05% 8.35 0.1056 0.0121 1.5344
Mistral-7B-v0.3 (Few-shot) 39.77% 38.92% 35.86% 34.31% 7.55 0.1144 0.0152 0.3023
Meta-Llama-3.1-8B (Zero-shot) 39.38% 32.65% 31.99% 29.45% 8.59 0.1068 0.0078 1.5218
Meta-Llama-3.1-8B (Few-shot) 39.96% 38.26% 35.77% 34.91% 7.70 0.1185 0.0139 0.4470
Qwen2.5-7B (Zero-shot) 36.88% 38.20% 37.69% 29.78% 9.50 0.1040 0.0123 1.7000
Qwen2.5-7B (Few-shot) 44.69% 39.52% 38.51% 34.87% 7.60 0.1183 0.0148 0.3720
$SurveyLM_{\text{Mistral-7B-v0.3}}$ 57.13% 54.44% 53.36% 53.85% 7.64 0.1308 0.0249 0.2974
$SurveyLM_{\text{Meta-Llama-3\_1-8B}}$ 55.09% 54.77% 51.02% 52.14% 7.86 0.1341 0.0189 0.4583
$SurveyLM_{\text{Qwen2.5-7B}}$ 56.83% 42.74% 41.49% 41.52% 7.69 0.1281 0.0172 0.3851
Table 15: Task3 Attitude Stance Modeling Comprehensive Result.
B.4 Task4 Structured Response Modeling
[…]
Model ACC $\uparrow$ PRE $\uparrow$ RECALL $\uparrow$ F1 $\uparrow$ WD $\downarrow$
GPT-4o (Zero-shot) 46.56 37.68 23.94 22.05 2.1336
GPT-4o (Few-shot) 47.91 38.40 23.40 23.10 1.7895
Claude 3.7 Sonnet (Zero-shot) 42.11 21.98 22.91 18.44 1.7789
Claude 3.7 Sonnet (Few-shot) 42.95 22.30 22.40 19.50 1.6542
DeepSeek-R1 (Zero-shot) 47.21 9.22 7.54 7.11 1.9987
DeepSeek-R1 (Few-shot) 47.88 9.05 8.20 7.65 1.8527
R1-Distill-Qwen-14B (Zero-shot) 49.54 28.82 31.96 25.90 1.9295
R1-Distill-Qwen-14B (Few-shot) 50.29 29.70 31.60 26.80 1.7164
Qwen2.5-72B (Zero-shot) 42.20 5.24 3.87 3.80 1.8654
Qwen2.5-72B (Few-shot) 43.06 5.60 3.70 4.05 1.7348
Meta-Llama-3.1-8B (Zero-shot) 14.84 0.37 0.13 0.17 2.1154
Meta-Llama-3.1-8B (Few-shot) 17.26 0.42 0.12 0.19 1.9451
Mistral-7B-v0.3 (Zero-shot) 28.04 1.85 1.07 1.10 2.2118
Mistral-7B-v0.3 (Few-shot) 30.47 2.10 0.95 1.25 2.0823
Qwen2.5-7B (Zero-shot) 16.39 0.27 0.10 0.12 1.8479
Qwen2.5-7B (Few-shot) 18.46 0.24 0.14 0.15 1.7059
$AlignSurvey_{\text{Mistral-7B-v0.3}}$ 54.59 45.30 41.01 41.34 0.0671
$AlignSurvey_{\text{Meta-Llama-3.1-8B}}$ 56.11 44.97 41.68 41.47 0.0675
$AlignSurvey_{\text{Qwen2.5-7B}}$ 55.47 44.23 41.87 41.79 0.0541
Table 16: Task4 Structured Response Modeling Results on the ASE.
Model ACC $\uparrow$ PRE $\uparrow$ RECALL $\uparrow$ F1 $\uparrow$ WD $\downarrow$
GPT-4o (Zero-shot) 53.43 32.18 30.61 26.24 1.9325
GPT-4o (Few-shot) 54.32 32.74 30.20 27.30 1.6992
Claude 3.7 Sonnet (Zero-shot) 47.21 12.50 10.54 9.50 1.8872
Claude 3.7 Sonnet (Few-shot) 47.98 13.02 11.20 9.95 1.7318
DeepSeek-R1 (Zero-shot) 44.65 4.55 8.65 4.52 1.7789
DeepSeek-R1 (Few-shot) 45.27 4.83 9.10 4.90 1.6523
R1-Distill-Qwen-14B (Zero-shot) 50.62 31.77 37.76 29.36 1.9448
R1-Distill-Qwen-14B (Few-shot) 51.37 32.35 37.10 30.67 1.7104
Qwen2.5-72B (Zero-shot) 47.98 18.23 18.02 13.29 1.9918
Qwen2.5-72B (Few-shot) 48.66 18.58 17.80 14.02 1.7989
Meta-Llama-3.1-8B (Zero-shot) 20.84 4.30 2.09 2.40 1.7748
Meta-Llama-3.1-8B (Few-shot) 22.46 4.68 1.95 2.60 1.6435
Mistral-7B-v0.3 (Zero-shot) 23.60 5.73 3.63 3.46 1.9941
Mistral-7B-v0.3 (Few-shot) 26.89 6.12 3.40 3.70 1.8876
Qwen2.5-7B (Zero-shot) 51.95 13.44 17.05 12.51 1.9099
Qwen2.5-7B (Few-shot) 53.14 13.98 17.55 13.15 1.7441
$AlignSurvey_{\text{Mistral-7B-v0.3}}$ 65.56 35.90 38.49 35.35 0.0857
$AlignSurvey_{\text{Meta-Llama-3.1-8B}}$ 64.89 32.91 36.82 32.99 0.0845
$AlignSurvey_{\text{Qwen2.5-7B}}$ 67.17 36.48 38.82 34.94 0.0686
Table 17: Task4 Structured Response Modeling Results on the CHIP.
Model ACC $\uparrow$ PRE $\uparrow$ RECALL $\uparrow$ F1 $\uparrow$ WD $\downarrow$
GPT-4o (Zero-shot) 32.79 27.19 26.94 24.41 1.1130
GPT-4o (Few-shot) 33.85 27.90 26.40 25.10 1.0284
Claude 3.7 Sonnet (Zero-shot) 33.94 30.72 28.41 25.93 1.2248
Claude 3.7 Sonnet (Few-shot) 34.55 30.50 28.10 26.60 1.1095
DeepSeek-R1 (Zero-shot) 34.54 29.64 27.63 23.15 1.3854
DeepSeek-R1 (Few-shot) 35.20 29.20 28.00 23.70 1.2432
R1-Distill-Qwen-14B (Zero-shot) 29.39 13.13 16.06 11.57 0.7920
R1-Distill-Qwen-14B (Few-shot) 30.20 13.60 15.80 12.25 0.7548
Qwen2.5-72B (Zero-shot) 34.50 22.93 25.75 21.14 1.2180
Qwen2.5-72B (Few-shot) 35.60 23.60 25.20 22.05 1.1112
Meta-Llama-3.1-8B (Zero-shot) 1.61 0.25 0.05 0.08 1.5004
Meta-Llama-3.1-8B (Few-shot) 6.90 0.40 0.04 0.12 1.3817
Mistral-7B-v0.3 (Zero-shot) 3.19 0.78 0.14 0.20 1.1479
Mistral-7B-v0.3 (Few-shot) 10.30 1.05 0.12 0.24 1.0913
Qwen2.5-7B (Zero-shot) 25.33 5.25 5.95 4.30 1.2411
Qwen2.5-7B (Few-shot) 26.85 5.10 6.30 4.55 1.1371
$AlignSurvey_{\text{Mistral-7B-v0.3}}$ 51.06 38.50 36.39 35.73 0.2097
$AlignSurvey_{\text{Meta-Llama-3.1-8B}}$ 51.44 35.01 34.66 33.84 0.2137
$AlignSurvey_{\text{Qwen2.5-7B}}$ 51.52 37.55 36.79 36.22 0.1849
Table 18: Task4 Structured Response Modeling Results on the GSS.
Across all tasks, SurveyLM models demonstrate clear superiority over general-purpose LLMs, highlighting the importance and effectiveness of domain-specific fine-tuning using the AlignSurvey benchmark and dataset. The consistent and significant improvements emphasize the necessity of specialized model training to achieve robust and socially representative alignment in modeling human preferences and attitudes across diverse survey tasks.
C. Extended Analysis
C.1 Ablation Studies
Task 1: Social Role Modeling.
Table 19 provides detailed ablation results for Task 1, assessing the impact of foundational training and task-specific supervised fine-tuning (SFT).
[…]
Model Basic INFO Social INFO Family INFO
Gender Age Area Edu. Employ. Stress Home Own. Veh. Own. Income Rel. HH Size Expenses
$SurveyLM_{\text{Mistral-7B-v0.3}}$ 54.31% 45.87% 59.67% 37.90% 43.47% 35.44% 60.54% 60.71% 33.18% 85.88% 39.99% 29.28%
   w/o Foundation 52.92% 45.42% 58.75% 37.08% 41.67% 33.75% 59.17% 59.58% 32.08% 85.42% 39.17% 27.92%
   w/o Task SFT 37.96% 28.32% 23.53% 33.92% 37.51% 30.39% 49.45% 56.77% 29.61% 76.21% 35.95% 16.38%
$SurveyLM_{\text{Meta-Llama-3\_1-8B}}$ 55.14% 47.03% 62.05% 41.02% 47.49% 44.02% 65.92% 65.01% 31.29% 80.06% 34.80% 32.78%
   w/o Foundation 53.75% 46.25% 60.42% 39.58% 46.67% 42.50% 64.58% 64.17% 30.42% 78.33% 33.75% 32.08%
   w/o Task SFT 39.64% 40.14% 57.69% 38.32% 44.66% 27.35% 60.94% 28.22% 18.90% 55.52% 28.99% 27.63%
$SurveyLM_{\text{Qwen2.5-7B}}$ 67.81% 50.62% 60.58% 41.83% 52.43% 36.55% 68.25% 66.88% 40.42% 86.42% 34.04% 33.95%
   w/o Foundation 66.25% 49.58% 60.00% 41.25% 50.42% 35.00% 67.08% 65.42% 39.58% 84.17% 32.92% 32.50%
   w/o Task SFT 41.87% 37.83% 36.72% 33.83% 13.51% 18.40% 52.96% 55.29% 38.79% 41.42% 29.07% 20.31%
Table 19: Ablation results for Task 1 under full model, without foundational training, and without task‐specific SFT. Higher percentages indicate better performance.
Task 2: Semi-Structured Interview Modeling.
Table 20 illustrates ablation effects on interview response generation, focusing on ”Naturalness,” ”Style Match,” and ”Consistency.”
[…]
Model Naturalness Style Match Consistency
$SurveyLM_{\text{Mistral}}$ 2.61 2.90 2.90
   w/o Foundation 3.5 2.90 2.90
   w/o Task SFT 3.08 2.90 2.90
$SurveyLM_{\text{Llama}}$ 3.77 3.00 2.94
   w/o Foundation 3.41 2.90 2.89
   w/o Task SFT 3.21 2.90 2.91
$SurveyLM_{\text{Qwen}}$ 3.98 2.96 3.01
   w/o Foundation 3.51 2.91 2.95
   w/o Task SFT 3.15 2.52 2.69
Table 20: Ablation study for Task 2 Semi-Structured Interview Modeling. Scores (1 = worst, 5 = best) are reported for Naturalness, Style Match, and Consistency on three SurveyLM backbones together with two variants of each: w/o Foundation (foundation pre-training removed) and w/o Task SFT (task-specific fine-tuning removed).
Task 3: Attitude Stance Modeling Table 21 evaluates ablation results on various stance metrics, including Accuracy, Precision, Recall, F1, and Wasserstein Distance (WD).
[…]
Model ACC PRE RECALL F1 ROUGE Jaccard Cosine WD
$AlignSurvey_{\text{Mistral}}$ 57.13% 54.44% 53.36% 53.85% 7.64 0.1308 0.0249 0.2974
   w/o Foundation 56.25% 53.57% 52.84% 53.01% 7.58 0.1292 0.0240 0.3026
   w/o Task SFT 43.75% 0.3871 0.3856 0.3349 6.82 0.1126 0.0101 0.3130
$AlignSurvey_{\text{Llama}}$ 55.09% 54.77% 51.02% 52.14% 7.86 0.1341 0.0189 0.4583
   w/o Foundation 54.37% 54.02% 50.14% 51.23% 7.75 0.1329 0.0182 0.4625
   w/o Task SFT 36.25% 0.3102 0.3262 0.2573 5.54 0.0940 0.0019 0.4840
$AlignSurvey_{\text{Qwen}}$ 56.83% 42.74% 41.49% 41.52% 7.69 0.1281 0.0172 0.3851
   w/o Foundation 56.25% 41.15% 40.91% 40.95% 7.65 0.1269 0.0165 0.3888
   w/o Task SFT 46.88% 61.09% 28.25% 21.94% 8.84 0.1110 0.0066 0.3970
Table 21: Ablation study for Task 3: Alignment Verification. We report classification metrics plus text-similarity metrics ROUGE-L, Jaccard, cosine similarity of sentence embeddings (higher is better), and WD (lower is better).
Task 4: Structured Response Modeling
Tables 22, 23, and 24 present ablation outcomes for structured survey response tasks across ASE, CHIP, and GSS datasets.
[…]
Model ACC PRE RECALL F1 WD
$AlignSurvey_{\text{Mistral}}$ 54.59% 45.30% 41.01% 41.34% 0.0671
   w/o Foundation 53.22% 44.12% 39.81% 39.57% 0.0765
   w/o Task SFT 36.17% 23.94% 23.89% 21.13% 2.182
$AlignSurvey_{\text{Llama}}$ 56.11% 44.97% 41.68% 41.47% 0.0675
   w/o Foundation 54.59% 44.29% 39.12% 38.50% 0.0846
   w/o Task SFT 34.33% 22.58% 19.70% 18.43% 2.098
$AlignSurvey_{\text{Qwen}}$ 55.47% 44.23% 41.87% 41.79% 0.0541
   w/o Foundation 54.86% 44.10% 39.55% 38.71% 0.8893
   w/o Task SFT 35.50% 21.63% 18.62% 17.52% 2.002
Table 22: Ablation results on ASE.
Model ACC PRE RECALL F1 WD
$AlignSurvey_{\text{Mistral}}$ 65.56% 35.90% 38.49% 35.35% 0.0857
   w/o Foundation 64.55% 33.87% 36.10% 32.54% 0.0912
   w/o Task SFT 40.32% 34.31% 33.55% 29.72% 2.000
$AlignSurvey_{\text{Llama}}$ 64.89% 32.91% 36.82% 32.99% 0.0845
   w/o Foundation 64.55% 34.64% 36.76% 33.61% 0.0900
   w/o Task SFT 40.03% 28.91% 29.45% 23.64% 1.984
$AlignSurvey_{\text{Qwen}}$ 67.17% 36.48% 38.82% 34.94% 0.0686
   w/o Foundation 64.07% 32.44% 36.30% 33.06% 0.0782
   w/o Task SFT 30.87% 19.18% 19.16% 15.12% 1.992
Table 23: Ablation results on CHIP.
Model ACC PRE RECALL F1 WD
$AlignSurvey_{\text{Mistral}}$ 51.06% 38.50% 36.39% 35.73% 0.2097
   w/o Foundation 50.34% 34.43% 34.88% 33.73% 0.3205
   w/o Task SFT 50.97% 37.10% 35.34% 34.56% 2.008
$AlignSurvey_{\text{Llama}}$ 51.44% 35.01% 34.66% 33.84% 0.2137
   w/o Foundation 51.48% 35.02% 34.45% 33.37% 0.2209
   w/o Task SFT 50.34% 34.60% 34.13% 33.05% 1.977
$AlignSurvey_{\text{Qwen}}$ 51.52% 37.55% 36.79% 36.22% 0.1849
   w/o Foundation 51.63% 37.94% 36.13% 35.46% 0.2118
   w/o Task SFT 50.47% 35.65% 34.59% 33.67% 1.992
Table 24: Ablation results on GSS.
C.2 Analysis by Demographic Sub-groups
[…]
Figure 4: Multi-Demographic Accuracy Comparison in Task3. Circle size represents the accuracy gap between Qwen-7B, Mistral-7B and LLaMA-8B, with their SurveyLM. Gains are especially pronounced for under-represented groups (rural residents, aged 76+, self-employed, low- and middle- income groups).
[…]
These demographic subgroup analyses confirm SurveyLM models’ superior alignment with diverse populations, effectively reducing response biases and improving representation accuracy across varied social and economic backgrounds.
Figure 5: Multi-Demographic Accuracy Comparison in Task4. Circle size represents the accuracy gap between Qwen-7B, Mistral-7B and LLaMA-8B, with their SurveyLM.
Extended analyses through ablation and demographic subgroups robustly confirm the essential role of task-specific fine-tuning and foundational training for achieving high-quality, fair, and demographically representative outcomes. SurveyLM models demonstrate remarkable improvements across demographic diversity, validating the significance of targeted model adaptation in enhancing both performance and fairness in social survey modeling.
D. Task Prompt Templates
D.1 Task 1 Social Role Modeling
The prompt template provided in Table 25 for Task 1 focuses on social role modeling by leveraging dialogue-based interactions. The task simulates an analytical scenario where the model, acting as an interview analyst, predicts demographic information (in this case, age) based on conversational cues provided within an interview.
The instruction explicitly defines the role of the model as an analyst tasked with interpreting survey dialogue to predict demographic attributes. The dialogue excerpt provided serves as a realistic representation of how respondents typically articulate their understanding and perception of social issues. The prompt clearly defines a query, which in this scenario is the respondent’s age, providing discrete age ranges as potential outcomes.
INSTRUCTION: You are an interview questionnaire analyst, analyzing a survey interview text about people’s perceived understanding of common prosperity. ==Interview Dialogue== Interviewer: Okay, then, in your mind, what should common prosperity look like? Interviewee: I believe common prosperity has two components. First, it’s universal prosperity, meaning it encompasses all people, not just a few. Second, I believe common prosperity is comprehensive. Beyond a sense of material prosperity, it also requires spiritual enrichment.
QUERY: Based on the conversation, please predict the respondent’s age.
Optional values: 18–25, 26–35, 36–45, 46–55, 56–65, 66–75, 76 and older.
ANSWER: 26–35 years old
Table 25: The Prompt Temple of Task1.
D.2 Task 2 Semi-Structured Interview Modeling
Task 2, illustrated in Table 26, represents a semi-structured interview modeling task. Here, the prompt situates the model as a survey respondent, characterized by specific demographic and socioeconomic attributes provided in detail within the background information section. The model is expected to produce responses consistent with this detailed profile, thereby assessing its ability to simulate realistic, context-aware human-like conversational behaviors.
The instructions delineate the scenario clearly, including previous conversational context, which the model must consider to ensure coherence and realism. The current question ”How would you put it?” prompts the model to generate a response consistent with the character’s socioeconomic status, education level, living situation, and employment status, reflecting nuanced awareness of subtle social distinctions.
INSTRUCTION: You are a respondent conducting a survey on the public’s perceived understanding of common prosperity. ==Your Background== Gender: Male Age: 26–35 Place of Household Registration: Taizhou, Zhejiang Province Permanent Residence: Hangzhou, Zhejiang Province Household Registration Status: Rural Highest Education Level: Graduate or above Number of Family Members: 3 Family Membership: Children, Grandparents Property Status: Yes Vehicle Status: No Annual Household Income: Over 1 million Major Monthly Household Expenditure Items: Daily Living Expenses (food, clothing, etc.); Education Expenses (tuition, supplementary materials, etc.); Healthcare Expenses (medical insurance, medicine, medical treatment, etc.) Life Stress Level: Mild Stress – occasionally feel stressed, but it does not affect life Occupation Type: Corporate Employee ==Previous Conversation== Question: Zhejiang began developing the Common Prosperity Demonstration Zone in 2021. Do you think there have been any changes compared to the past, especially before the establishment of the Common Prosperity Demonstration Zone? Answer: I personally think the difference is not significant. ==Current Question== How would you put it?
QUERY: Please enter your answer:
ANSWER: Maybe it is related to the industry I work in. I work in the Internet industry. Relatively speaking, the overall salary of the industry and the company has not changed much in the past two years.
Table 26: The Prompt Temple of Task2.
D.3 Task 3 Attitude Stance Modeling
The prompts in Tables 27 and 28 outline two critical subtasks within attitude stance modeling. Task 3.1 requires the model to analytically determine and rationalize the respondent’s stance towards public services based on demographic data and implied context. The instruction specifies an output format, demanding structured, interpretable reasoning that explicitly justifies the model’s stance prediction in JSON format.
Task 3.2 is structured quantitatively, requiring the model to predict response distribution across positive, neutral, and negative attitudes toward common prosperity among a specified demographic subgroup (26-35 years). The prompt explicitly demands a percentage-based distribution output, emphasizing the model’s capability to aggregate nuanced attitude information into accurate, representative quantitative summaries.
INSTRUCTION: You are a sociology researcher analyzing interview data from a respondent: ==Your Background== Gender: Male Age: 26–35 Place of Household Registration: Taizhou, Zhejiang Province Permanent Residence: Hangzhou, Zhejiang Province Household Registration Status: Rural Highest Education Level: Graduate or above Number of Family Members: 3 Family Membership: Children, Grandparents Property Status: Yes Vehicle Status: No Annual Household Income: Over 1 million Major Monthly Household Expenditure Items: Daily Living Expenses (food, clothing, etc.); Education Expenses (tuition, supplementary materials, etc.); Healthcare Expenses (medical insurance, medicine, medical treatment, etc.) Life Stress Level: Mild Stress – occasionally feel stressed, but it does not affect life Occupation Type: Corporate Employee
QUERY: Please determine the respondent’s evaluation of public services and explain your reasoning.
The output format is JSON and contains two fields: ”predict” and ”reason.” The predict field is selected from ”positive”, ”negative” or ”neutral”.
ANSWER: {”predict”: ”Positive”, ”reason”: ”Respondents have lived in county towns for a long time and may have been influenced by the renovation of older residential areas and government public service improvement projects. Therefore, they have a relatively positive attitude towards public services, believing them to have ”improved somewhat.” Regarding education and healthcare, respondents expressed limited understanding. This may be due to the health of their elderly family members and their current childlessness, which may result in relatively low demand for education and healthcare, and limited experience or feelings about them.”}
Table 27: The Prompt Temple of Task3.1.
INSTRUCTION: Please provide the distribution of responses to the following question for people aged 26-35: What is your attitude towards common prosperity? 1. Positive 2. Neutral 3. Negative
QUERY: Please provide the distribution of responses (in percentages), for example,
A.xx: xx%
B.xx: xx%
C.xx: xx%
ANSWER: {”Positive”: ”40%”, ”Neutral”: ”46%”, ”Negative”: ”13%”}
Table 28: The Prompt Temple of Task3.2.
D.4 Task 4 Structured Response Modeling
The structured response modeling task, described in Tables 29 and 30, focuses on both individual and aggregate structured responses to survey questions. Task 4.1 simulates a realistic scenario wherein the model represents a respondent with explicitly provided demographic and socioeconomic contexts. This explicit profile requires the model to generate responses that accurately reflect likely perspectives based on specified individual attributes, such as place of residence, income, family structure, and occupation.
Task 4.2 extends structured response modeling by requiring demographic-specific distribution predictions for community engagement attitudes among office-working females with moderate household incomes. This task demands sophisticated aggregation capabilities, as the model must synthesize demographic knowledge to generate believable statistical distributions. The explicitly percentage-based format tests both demographic sensitivity and the model’s ability to accurately capture population-level variability and representativeness in responses.
INSTRUCTION: Survey time: 2025. You need to simulate a real survey taker to answer the following questions. ==Your Background== Gender: Male Age: 26–35 Place of Household Registration: Taizhou, Zhejiang Province Permanent Residence: Hangzhou, Zhejiang Province Household Registration Status: Rural Highest Education Level: Graduate or above Number of Family Members: 3 Family Membership: Children, Grandparents Property Status: Yes Vehicle Status: No Annual Household Income: Over 1 million Major Monthly Household Expenditure Items: Daily Living Expenses (food, clothing, etc.); Education Expenses (tuition, supplementary materials, etc.); Healthcare Expenses (medical insurance, medicine, medical treatment, etc.) Life Stress Level: Mild Stress – occasionally feel stressed, but it does not affect life Occupation Type: Corporate Employee
QUERY: To what extent do you believe the economy of your county (city, district) is developing rapidly?
Options:
A. Neutral
B. Agree
C. Disagree
Please enter your answer directly.
ANSWER: B. Agree
Table 29: The Prompt Temple of Task4.1.
INSTRUCTION: Please provide the distribution of the following groups of people across the questionnaire options below. ==Population Characteristics== Gender: Female Occupation: Office Worker Annual Household Income: 110,000-200,000 RMB
QUERY: To what extent do you believe you can participate in community or village affairs?
Options:
A. Agree
B. Neutral
C. Disagree
Please provide the distribution of answers directly (in percentage form), for example:
A. xx%
B. xx%
C. xx%
ANSWER:
Neutral: 47%
Disagree: 16%
Agree: 36%
Table 30: The Prompt Temple of Task4.2.
By encompassing diverse survey response formats—from open-ended dialogues to structured quantitative responses—the provided templates offer versatile benchmarks essential for developing and refining socially aligned language modeling technologies. Consequently, they serve as foundational tools for systematically advancing model interpretability, accuracy, realism, and social representativeness in human-oriented AI applications.
E. Automatic LLM-based Evaluators
E.1 Evaluator Prompt Template
The prompt used for the automatic evaluators is shown in Table 31.
E.2 Scoring Dimensions and Rating Criteria
The evaluator scores each mock answer along three dimensions: Naturalness, Style Match, and Consistency. Each dimension is rated on a 1 – 5 ordinal scale.
INSTRUCTION: Mock Interview Response Evaluation Criteria Scoring Dimensions 1. Naturalness: Evaluate whether the interview response is expressed in natural language, connects well with the questions, and reflects the tone and rhythm of real-life conversation. Avoid templated or bulleted responses. **If the response is too structured, the score for this item should be appropriately reduced. 2. Style Match: Evaluate whether the response is consistent with the standard response, including language style, thinking style, and focus. 3. Consistency: Determine whether the response accurately expresses the core points related to the question and covers the main positions or logic reflected in the standard answer. **If the response expresses judgments not reflected in the standard answer, the score for this item should be appropriately reduced. Rating Criteria Score the mock answers based on the above dimensions, using a 1-5 scale for each dimension. - 5: Fully meets the evaluation criteria, with accurate, natural, and complete expression. - 4: Mostly meets the criteria, with minor deficiencies or deviations. - 3: Partially meets the criteria, with significant problems or omissions. - 2: Only partially meets the criteria, with numerous problems that hinder comprehension. - 1: Completely fails to meet the criteria, with disorganized structure or significant deviation from the topic. ==Your Background== Gender: Male Age: 26–35 Place of Household Registration: Taizhou, Zhejiang Province Permanent Residence: Hangzhou, Zhejiang Province Household Registration Status: Rural Highest Education Level: Graduate or above Number of Family Members: 3 Family Membership: Children, Grandparents Property Status: Yes Vehicle Status: No Annual Household Income: Over 1 million Major Monthly Household Expenditure Items: Daily Living Expenses (food, clothing, etc.); Education Expenses (tuition, supplementary materials, etc.); Healthcare Expenses (medical insurance, medicine, medical treatment, etc.) Life Stress Level: Mild Stress – occasionally feel stressed, but it does not affect life Occupation Type: Corporate Employee ==Previous Conversation== Question: Zhejiang began developing the Common Prosperity Demonstration Zone in 2021. Do you think there have been any changes compared to the past, especially before the establishment of the Common Prosperity Demonstration Zone? Answer: I personally think the difference is not significant. ==Current Question== How would you put it? Model Answer Maybe it is related to the industry I work in. I work in the Internet industry. Relatively speaking, the overall salary of the industry and the company has not changed much in the past two years. Mock Answer Being part of the Internet industry might explain it. Over the past two years, compensation for both the sector as a whole and my employer in particular has remained largely unchanged.
QUERY: Output Format {{ ”Naturalness”: {{ ”Rating”: ”X”, ”Comment”: ”…” }}, ”Style Match”: {{ ”Rating”: ”Y”, ”Comment”: ”…” }}, ”Consistency”: {{ ”Rating”: ”Z”, ”Comment”: ”…” }} }}
ANSWER: {”Naturalness”: { ”Rating”: ”4”, ”Comment”: ”The simulated answers were expressed in natural language and connected to the questions reasonably, but lacked specific examples, making the conversation a bit vague.}, ”Style Match”: { ”Rating”: ”3”, ”Comment”: ”The focus of the answer deviated from the specific examples in the standard answer, while the language style remained consistent but slightly vague.}, ”Consistency”: {”Rating”: ”3”, ”Comment”: ”The specific examples in the standard answer were inconsistent, failing to accurately express the specific perception of public service use and lacking specific examples.}
Table 31: The Prompt Temple of Evaluator in Task2.
E.3 Alignment Evaluation
Statistic Human–Human Human–LLM
Fleiss’ $\kappa$ 0.68
Krippendorff’s $\alpha$ 0.72
Pearson $r$ 0.74
Spearman $\rho$ 0.73
Table 32: Inter-annotator reliability and human–model alignment statistics. Values above 0.60 (reliability) and 0.70 (correlation) are generally considered acceptable.
To verify the reliability of the automatic LLM scores, we first collected pairwise comparative judgments from domain experts. For each prompt, three experts compared two model responses on overall preference. The consistency of our annotation rubric was assessed with both Fleiss’ $\kappa$ and Krippendorff’s $\alpha$. Both statistics slightly exceed the commonly accepted 0.60 threshold, indicating moderate but acceptable inter-rater agreement.
We then examined the alignment between aggregated human scores and those produced by an automatic LLM-based evaluator (Table 32). Pearson and Spearman correlations in the higher than 0.70 show that the evaluator captures human preferences reasonably well while providing orders-of-magnitude higher throughput.
F. Annotation Reliability
To operationalise Task 3 Attitude Stance Modeling, we first created a concise codebook that maps every dialogue segment to one of three mutually–exclusive stance labels: positive (1), neutral (2), or negative (3). Three graduate-level social-science researchers were recruited. Throughout production coding, annotators worked independently and blind to one another.
Every dialogue segment (N = 644) received three independent labels. Disagreements triggered a two-step adjudication:
  • Automatic majority rule: if two of three raters agreed, their label became the provisional gold.
  • Expert arbitration: segments with a three–way split were forwarded to a senior domain expert, who reviewed the raw transcript together with the annotators’ written rationales and issued a final decision.
We computed three complementary reliability statistics:
Statistic Rater(s) Compared $\kappa$ / $\alpha$ $P_{o}$
Cohen’s $\kappa$ people1 vs. people2 0.63 0.793
people1 vs. people3 0.71 0.852
people2 vs. people3 0.66 0.811
Fleiss’ $\kappa$ all three 0.67 0.822
Krippendorff’s $\alpha$ all three 0.74
Table 33: Inter-coder agreement statistics for Task 3.
Taken together, these scores show that annotators converged well above chance and satisfy the minimum thresholds ($\geq$ 0.60 for $\kappa$, $\geq$ 0.67 for $\alpha$) recommended for corpus construction in empirical social-science research.
Although the stance scale is deliberately fine-grained, moderate–to–substantial agreement was achieved without prolonged training, suggesting that the codebook captures readily observable cues in ordinary language. The residual disagreements mainly stem from (i) sarcastic or ambivalent statements and (ii) domain-specific jargon. Future work can further boost consistency by providing additional sarcasm exemplars and a mini-glossary for specialised terms.
G. Licensing, Ethics & Societal Impact
G.1 Data Licenses and Usage Rights
All datasets used in the AlignSurvey project were sourced from publicly accessible platforms or officially released academic resources, and processed solely for non-commercial research purposes in accordance with their respective usage terms and academic licensing policies.
Video and Interview Text Data: The data were collected from publicly accessible platforms and oral history materials, restricted to content openly available without access limitations. All processing was conducted for academic use only. The data construction process involved transcription and anonymization of spoken content. No original video, audio, or full dialogue content is stored, retained, or distributed. All downstream resources are derived exclusively from transcribed text and expert annotations. Prior peer-reviewed research has demonstrated the value and academic acceptability of using content from public video platforms for large-scale benchmarking and social analysis (real2017youtube; loh2022youtube; albadi2022deradicalizing). Our approach is consistent with this line of work and adheres to public content usage norms and academic best practices.
Social Survey Data: We incorporate publicly released, academically licensed survey datasets including the American Trends Panel (ATP), European Social Survey (ESS), General Social Survey (GSS), Chinese Social Survey (CSS), Chinese General Social Survey (CGSS), and China Household Income Project (CHIP). All datasets are used in strict accordance with their official licensing terms and intended for non-commercial academic research only. Personally identifiable information was thoroughly removed following established anonymization procedures to ensure participant privacy.
Data Distribution and Public Release: AlignSurvey-related data, tools, and model parameters are publicly released via platforms such as GitHub and HuggingFace under the Creative Commons Attribution–NonCommercial–NoDerivatives 4.0 International License (CC BY-NC-ND 4.0). All resources are intended strictly for non-commercial academic research. Users must strictly adhere to both the open-source license terms and the licensing agreements of any original data sources included in the benchmark.
G.2 Model Release and License Compliance
The SurveyLM model family released by AlignSurvey is based on open-source large language models (LLMs), refined through a two-stage fine-tuning process. Initially, models are trained on a diverse social-context corpus, encompassing multilingual dialogues and structured survey texts to develop broad social-contextual awareness. Subsequently, these models undergo task-specific supervised fine-tuning with comprehensive datasets tailored to each research task. All model checkpoints are released strictly for non-commercial academic research purposes.
Mistral 7B Model: Released under the Apache 2.0 License, which allows broad usage—including modification and redistribution—provided that proper attribution and copyright notices are retained.
Meta Llama 3.1 Model: Released under Meta’s COMMUNITY LICENSE AGREEMENT. This license permits non-commercial research use and derivative works, while prohibiting commercial deployment without separate authorization.
Qwen-7B Model: Also released under the Apache 2.0 License, permitting academic and commercial use with proper attribution, in accordance with the license terms published by Qwen’s maintainers.
While the base models support varying degrees of commercial use, all SurveyLM models released as part of the AlignSurvey benchmark are explicitly scoped for non-commercial academic use only. Distributed model weights include clear notices reinforcing original license terms, derivative usage conditions, and the research-only scope of redistribution.
G.3 Ethical Compliance and Privacy Protection
AlignSurvey adheres to rigorous ethical and privacy standards throughout data curation and model training:
Privacy by Design: All interview and survey content underwent strict de-identification and removal of any potentially sensitive information. Dataset construction avoided collecting or storing any original media content (e.g., video, audio, or images).
Minimal Use and Responsible Processing: For publicly sourced materials, only transcribed and anonymized interview text was retained. No raw media content, full transcripts, or personally revealing information is included or shared. This aligns with academic norms and ensures the minimal-use principle is upheld.
Academic Usage Boundaries: All released resources are restricted to non-commercial academic research. They are not intended for use in clinical, legal, or other high-stakes decision-making scenarios without appropriate domain expertise and human oversight.
Inclusivity and Fairness: Throughout dataset construction and model evaluation, efforts were made to ensure broad demographic representation and minimize systematic bias. Particular care was given to underrepresented or marginalized groups, in both sampling and evaluation criteria.
G.4 Responsible Use and Ethical Safeguards
To support responsible innovation and uphold high standards of academic integrity, we implement the following safeguards:
Content Quality Assurance: Prior to release, model outputs were reviewed through a combination of automated tools and human-in-the-loop checks to ensure thematic relevance and discursive appropriateness. Users are encouraged to verify fitness for their specific research scenarios in accordance with institutional practices.
Usage Guidelines: The project provides clear documentation regarding usage scope, licensing, and intended applications. Users are expected to follow institutional research guidelines and uphold responsible academic practices.
These safeguards are designed to foster constructive use of large language models in computational social science while promoting transparency, inclusivity, and scholarly rigor.